I don't even know where to begin on this problem.
Cameron invested three times as much money at 7.95% as he invested at 6.5%. If his yearly simple interest is $282.26, how much money was invested at 6.5%?

I have a hard time solving these kinds of problems. I use a chart to work these problems out. It seems like so much is left out of the word problem.

Answer by LilSkittleMd(119)

There are 2 equations for this problems
0.065x+0.0795y=282.26
y=3x
Plug in the y= value into the equation
0.065x+0.0795(3x)=282.26
0.065x+0.2385x=282.26
0.3035x=282.26
x=930.02
y=3(930.02)
y=2790.06
$930.02 at 6.5%